Graph the function "f(x)=2x+5." Use the graph to find the indicated​ limit, if it exists.

"\lim\limits_{x\to\infin} f(x)".



1
Expert's answer
2021-05-27T12:22:12-0400

Let us graph the function "f(x)=2x+5:"





It follows that "\\lim\\limits_{x\\to\\infin} f(x)=\\lim\\limits_{x\\to\\infin} (2x+5)=\\infin.\n\n\u200b"